OpenStudy (anonymous):

Find two numbers whose product is a maximum when the sum of half the first number and ten times the second number is 100

OpenStudy (anonymous):

@gymnast4Jesus<3

OpenStudy (anonymous):

Oh, you are not so good in tagging someone.. :P

OpenStudy (anonymous):

haha i know im sorry

OpenStudy (anonymous):

do you have any idea how to solve this problem?

OpenStudy (anonymous):

@gorv wanna continue with Regina??

OpenStudy (anonymous):

please i beg. need all the help i can get

OpenStudy (gorv):

only if u r not willing to do it :P

OpenStudy (anonymous):

Okay, I am not willing.. :P

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

What is that simile for??

OpenStudy (gorv):

let first number is =x

OpenStudy (anonymous):

what is have so far ir 100=1/2x times 10z

OpenStudy (anonymous):

I *

OpenStudy (anonymous):

@Regina.D gorv will help you here.. :)

OpenStudy (anonymous):

Thank you gorv

OpenStudy (gorv):

second number =y

OpenStudy (anonymous):

okay im using x and y

OpenStudy (gorv):

\[\frac{ x }{ 2 }+10*y=100\]

OpenStudy (gorv):

subtarct x/2 both side

OpenStudy (anonymous):

okay i subtracted it

OpenStudy (gorv):

10y=100-x/2

OpenStudy (gorv):

divide both side by 10

OpenStudy (gorv):

y=10-x/20

OpenStudy (anonymous):

i wrote that

OpenStudy (gorv):

so we got the value of y

OpenStudy (gorv):

so now their product

OpenStudy (anonymous):

so now that we have the y value we need x right? since the problem asks for 2 numbers

OpenStudy (gorv):

no x we will keep and replace y with what we got

OpenStudy (anonymous):

oh

OpenStudy (gorv):

product= x*(10-x/20)

OpenStudy (gorv):

\[=10x-\frac{ x^2 }{ 20 }\]

OpenStudy (gorv):

we have to maximize it

OpenStudy (anonymous):

how do we maximize it?

OpenStudy (gorv):

using differentiation

OpenStudy (gorv):

on differentiation

OpenStudy (gorv):

=10-2*x/20

OpenStudy (anonymous):

oh god i dont know how to do that

OpenStudy (anonymous):

You can also maximize it by stretching it from both the sides.. :P

OpenStudy (anonymous):

my professor is an idiot

OpenStudy (gorv):

u know how to do differentiation

OpenStudy (anonymous):

i dont we werent taught it yet which is why im so confused with this problem

OpenStudy (anonymous):

This topic of maximize and minimize comes under Differentiation part...

OpenStudy (gorv):

okk..u know anything about differentiation

OpenStudy (anonymous):

nothing at all

OpenStudy (anonymous):

i just started pre calc

OpenStudy (gorv):

wait

OpenStudy (anonymous):

May be there is some other method without using differentiation.. :(

OpenStudy (anonymous):

that is why i'm here...because i need help

OpenStudy (gorv):

we can do it using quadratic 's graph but i know that ..but i had prob that i cant make u understand that

OpenStudy (gorv):

@ganeshie8 @hartnn plzz help here..

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

its not that hard of a problem..

OpenStudy (anonymous):

i;m just stupid in math

OpenStudy (anonymous):

but you guys will solve it for sure

OpenStudy (anonymous):

....................................

OpenStudy (anonymous):

alright you guys are all idiots lmao

OpenStudy (anonymous):

see ya

OpenStudy (anonymous):

We all are idiots here.. :) It is a site for idiots, that's why I think you came here.. :P

OpenStudy (gorv):

@amistre64

OpenStudy (anonymous):

May I try now??

OpenStudy (gorv):

loll if u know u can

OpenStudy (anonymous):

Let us wait and see what amistre has got in his pocket.. :)

OpenStudy (amistre64):

Find two numbers define the 2 numbers as: x and y whose product is a maximum xy is largest when ... when the sum of half the first number x/2 and ten times the second number 10y is 100 x/2 + 10y = 100

OpenStudy (anonymous):

Simply find x and y?? :P

OpenStudy (amistre64):

all i did was summaraize gorvs procees i believe

OpenStudy (gorv):

no after that..how we will maximise

OpenStudy (amistre64):

well, with substitution we get a quadratic ....

OpenStudy (gorv):

bcozz the asker havent done differentiation

OpenStudy (gorv):

yeah with quadratic we will do thta what i said

OpenStudy (amistre64):

a vertex is gone over in algebra

OpenStudy (gorv):

but i cant guide him/her so asked for your assistance

OpenStudy (gorv):

i m weak in qudratic graph having max and min prob

OpenStudy (anonymous):

Somewhere I read, that for quadratic equation's product to be maximum then it is maximum when : \(x = \frac{-b}{2a}\)

OpenStudy (amistre64):

x/2 + 10y = 100 solve for x (or y, im doing x) x = 200 - 20y sub into f = xy f = (200-20y)y f = 200y -20y^2 plenty of ways to find the vertex

OpenStudy (amistre64):

x and y in my case are just numbers, not necessarily axis names

OpenStudy (anonymous):

I mean when product is to maximum, then maximum occurs when : \[x = \frac{-b}{2a}\]

OpenStudy (amistre64):

what part of this is the asker having difficulty with ?

OpenStudy (anonymous):

For him or her, it is done upto Product.. after that............

OpenStudy (anonymous):

\[P = x y =10x-\frac{ x^2 }{ 20 }\]

OpenStudy (amistre64):

my x sub makes a prettier solution process :) but im biased

OpenStudy (anonymous):

Wait, after that what you want to do??

OpenStudy (anonymous):

You are going to find maxima for f and then all that?? Is that??

OpenStudy (anonymous):

By using Differentiation??